OpenStudy (anonymous):

dy/dx of x^3+y^3=4xy+1

hartnn (hartnn):

take it term by term, whats d/dx of x^3 =....? d/dx y^3 =...?? [Note : here you will need a chain rule]

OpenStudy (anonymous):

d/dx of x^3 = 3x^2 is the d/dx of y^3 not just 3x^2 as well? why do i need the chain rule?

hartnn (hartnn):

because y^3 is NOT the function of 'x' so, by chain rule, \(\Large \dfrac{d}{dx}y^3=3y^2\dfrac{dy}{dx}\) got this ?

OpenStudy (anonymous):

ohh, okay, yes, got that. i just didn't realize that was the chain rule, since i didn't see an "outer" or "inner" part.

hartnn (hartnn):

ok, so we have our left part, what about the right part ? d/dx [4xy] =..... ? [Note you will need product rule here.] d/dx 1 =... ?

OpenStudy (anonymous):

d/dx [4xy]= 4(x(dy/dx)+y) which = 4x(dy/dx)+4y ? d/dx 1 = 0

hartnn (hartnn):

correct! good :) so, we'll just need to isolate dy/dx from \(\Large 3x^2+3y^2\dfrac{dy}{dx}=4x\dfrac{dy}{dx}+4y\) try first ?

OpenStudy (anonymous):

dy/dx (3y^2-4x) = 4y-3x^2 ?

hartnn (hartnn):

oh my! you are good at this :D its correct :) just divide 3y^2-4x and you're done :)

OpenStudy (anonymous):

haha, thanks... i understand majority of implicit differentiation, i just need a little help on a few steps here and there. so my answer would be : dy/dx=4y-3x^2/3y^2-4x ? if so, then why does it come out to 3x^2-4y/4x-3y^2 instead on wolfram alpha?

hartnn (hartnn):

they both are same! (a-b)/(c-d) is same as (b-a)/(d-c)!

hartnn (hartnn):

both are perfectly valid and accurate answers :)

OpenStudy (anonymous):

oh, duh. i remember that now. well thank you so much!!

hartnn (hartnn):

you're most welcome ^_^ and since you are new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{e}}\color{#e6e600}{\text{l}}\color{green}{\text{c}}\color{blue}{\text{o}}\color{purple}{\text{m}}\color{purple}{\text{e}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{S}}\color{#e6e600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{ }}\color{#e6e600}{\text{:}}\color{green}{\text{)}}\color{blue}{\text{}}\end{array} \]

hartnn (hartnn):

if you have any doubts browsing this site, you can ask me :)

OpenStudy (anonymous):

sweet, thanks! if i come up with more questions (which i undoubtedly will), is there a way i can just ask you directly?

hartnn (hartnn):

you can message me, but a preferable choice will be to post it here,as a new question, so that even if i am not online, you can surely get help from some one else :)